Will solve the equation. Just need help utilizing proper formula. Thank you. How many gallons of a 3% salt solution must be mixed with 50 gallons of a 7% solution to obtain a 5% solution?
Answer by mananth(16949)

percent ----------------quantity
7 ----------------50
3 ----------------x
5 ----------------50+x
..
7*50+3*x=5(50+x)
350+3x=250+5x
3x-5x=250-350
-2x=-100
/-2
x=50 gallons